1. Add 0.05moll-mgcl 2 solution.Can precipitation occur when ammonia water is mixed in equal volume? to 0.1mol · L-
1
Expert's answer
2019-12-03T08:23:51-0500

Solution.

"MgCl2 + 2NH4OH = Mg(OH)2 + 2NH4Cl"

"Cnew(MgCl2) = \frac{C(MgCl2)}{2}"

"Cnew(NH4OH) = \frac{C(NH4OH)}{2}"

"Ksp = (0.025)\times(0.05)^2 = 6.25 \times 10^{-5}"

"Ksp(theor.) = 6.0\times10^{-10}"

Since the theoretical product of the solubility of magnesium hydroxide was less than the experimental one, the precipitate will fall out.

Answer:

Since the theoretical product of the solubility of magnesium hydroxide was less than the experimental one, the precipitate will fall out.

"Ksp = 6.25 \times 10^{-5}"